OpenStudy (anonymous):

If the mass of an apple is 100g on average and the mass of the earth is approximately 6.0 * 10^24kg how many apples would it take to equal the mass of the earth? How many moles of apple does this represent?

sam (.sam.):

an apple is 100g we convert it to 0.1kg for easier calculation Let n=number of apples ---------------------------- We want the mass of all the apples equal to the mass of the earth, \[n(0.1)=6.0 * 10^{24}\] \[n=6\times 10^{25}\] ===================== number of apples is \(6\times 10^{25}\)

sam (.sam.):

So, the mass of \(6\times 10^{25}\) apples is \(( 6\times 10^{25} )\times 100g\) \(=6 \times 10^{27}g\) \[6 \times 10^{27}g \times\frac{1~mole}{6.02 \times 10 ^{23}~g}=9966~moles\]

OpenStudy (anonymous):

Thx :D